Problem 3-15 Liquidity Ratios (LG3-1) makes up a r 40 percent of current assets. Brendas current ratio is 1.9 times Calculate the value of inventory listed on the firms balance sheet. (Enter your answer In millions of dollars of in
Answer #1

Current ratio=current assets/current liabilities

Hence current assets=(17*1.9)=$32.3 million

Current assets=cash+accounts receivable+inventory

Cash and accounts receivable make up =(20+40)=60% of current assets

Hence inventory is =(100-60)=40% of current assets

Hence value of inventory=(32.3*40%)

=$12.92 million.